OpenStudy (anonymous):

Solve for x: 3|x - 3| + 2 = 14

OpenStudy (ipwnbunnies):

Subtract the 2, divide by 3. Left with abs(x-3) = '??' You want to split this up so x-3 = positive and negative '??'. This is because a negative number in absolute value is still a positive number.

OpenStudy (ipwnbunnies):

So you'll actually get 2 answers for x.

OpenStudy (anonymous):

x = -1, x = 7? @iPwnBunnies

OpenStudy (anonymous):

3|x-3|=14-2=12 |x-3|=12/3=4 |x-3|=4 x-3=4 ......or x-3=-4 x=3+4....or x=-4+3 x=7 ...or ...x=-1